Oxygen (O2) has a molar mass of 32.0g/mol .

What is the root-mean-square speed of an oxygen molecule at a temperature of 297K ?

What is its average translational kinetic energy at that speed?
Answer #1

M = 0.032 kg/mol,
R = gas const = 8.314472 J/mole-K,
k = Boltzmann's const R/Avogadro = 1.380657E-23 J/molecule-K
vRMS = sqrt(3RT/M) = 483.574735692426 m/s
Mean KE = 3(kB)T/2 = 6.21292644788425E-21 J